40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 304 of 357  Not logged in ELOG logo
ID Date Author Type Category Subject
  2693   Mon Mar 22 10:07:30 2010 KojiSummaryElectronicsUPDH Box #17: Ready

For your reference: Voltage noise of LM7815/LM7915 (with no load)

Attachment 1: 15V_power_supply.pdf
15V_power_supply.pdf
  2692   Mon Mar 22 02:03:57 2010 ranaSummaryElectronicsUPDH Box #17: Ready

It took too long to get this box ready for action. I implemented all of the changes that I made on the previous one (#1437). In addition, since this one is to be used for phase locking, I also made it have a ~flat transfer function. With the Boost ON, the TF magnitude will go up like 1/f below ~1 kHz.

The main trouble that I had was with the -12V regulator. The output noise level was ~500 nV/rHz, but there was a large oscillation at its output at ~65 kHz. This was showing up in the output noise spectrum of U1 (the first op-amp after the mixer). Since the PSRR of the OP27 is only ~40 dB at such a high frequency, it is not strange to see the power supply noise showing up (the input referred noise of the OP27 is 3.5 nV/rHz, so any PS noise above ~350 nV/rHz becomes relavent).

I was able to tame this by putting a 10 uF tantalum cap on the output of the regulator. However, when I replaced the regulator with a LM7912 from the blue box, it showed an output noise that went up like 1/f below 50 kHz !! I replaced it a couple more times with no benefit. It seems that something on the board must now be damaged. I checked another of the UPDH boxes, and it has the same high frequency oscillation but not so much excess voltage noise. I found that removing the protection diode on the output of the regulator decreased the noise by a factor of ~2. I also tried replacing all of the 1 uF caps that are around the regulator. No luck.

Both of the +12 V regulators seem fine: normal noise levels of ~200 nV/rHz and no oscillations.

Its clear that the regulator is not functioning well and my only guess is that its a layout issue on the board or else there's a busted component somewhere that I can't find. In any case, it seems to be functioning now and can be used for the phase locking and PZT response measurements.

  2691   Sun Mar 21 21:02:39 2010 KojiUpdatePSLEOM waist size
You don't need a lengthy code for this. It is obvious that the spot size at the distance L is minimum when L =
zR, where zR is the Rayleigh range. That's all.

Then compare the spot size and the aperture size whether it is enough or not.

It is not your case, but if the damage is the matter, just escape to the large zR side. If that is not possible
because of the aperture size, your EOM is not adequate for your purpose.
  2690   Sun Mar 21 20:08:20 2010 kiwamu, ranaUpdatePSLEOM wasit size

We are going to set the waist size to 0.1 mm for the beam going through the triple resonant EOM on a new PSL setup.

When we were drawing a new PSL diagram, we just needed to know the waist size at the EOM in order to think about mode matching.

waist.png

This figure shows the relation between the waist size and the spot size at the aperture of the EOM.

The x-axis is the waist size, the y-axis is the spot size. It is clear that there is a big clearance at 0.1 mm waist size. This is good.

Also it is good because the waist size is much above the damage threshold of the EO crystal (assuming 1W input).

The attached file is the python code for making this plot.

Attachment 2: waist.py.zip
  2689   Sun Mar 21 19:25:29 2010 ranaUpdatePEMGuralp2 centered again

From this morning, now in calibrated units, and with the Güralp self noise spec from the Güralp manual.

Attachment 1: huddlez.png
huddlez.png
  2688   Sat Mar 20 18:34:19 2010 kiwamuSummaryElectronicsRE:advantege of our triple resonant EOM

Yes, I found it.

Their advantage is that their circuit is isolated at DC because of the input capacitor.

And it is interesting that the performance of the circuit in terms of gain is supposed to be roughly the same as our transformer configuration.

  2687   Fri Mar 19 23:03:41 2010 ranaUpdatePEMGuralp2 centered again

I went and double-checked and aligned the styrofoam cooler at ~5:00 UTC. It was fine, but we really need a better huddling box. Where's that granite anyway?

Here's the new Huddle Test output. This time I show the X-axis since there's some coherence now below 0.1 Hz.

You'll also notice that the Wiener filter is now beating the FD subtraction. This happened when I increased the # of taps to 8000. Looks like the noise keeps getting lower as I increase the number of taps, but this is really a kind of cheat if you think about it carefully.

Attachment 1: huddlez.png
huddlez.png
  2686   Fri Mar 19 21:15:33 2010 ranaUpdatePEMGuralp2 centered again

This trend of the last 200 days shows that GUR2 has been bad forever...until now anyways.

Attachment 1: Untitled.png
Untitled.png
  2685   Fri Mar 19 18:00:14 2010 jenneUpdatePEMGuralp2 centered again

[Jenne, Sanjit]

It looks like Steve used a GND-12V supply to power the Guralp through the little breakout box (the box is for checking the centering of the mass).  This is BAD.  The Guralps want +/- 12V.

We centered all of the channels on Gur2, and checked the channels on Gur1, so we'll see how they're feeling after a while.

  2684   Thu Mar 18 21:42:26 2010 KojiUpdateABSLPLL reconstructed

I checked the setup further more.

  • I replaced the PD from NewFocus 1GHz one to Thorlabs PDA255.
  • I macthed the power level of the each beam.

Now I have significant fraction of beating (30%) and have huge amplitude (~9dBm).
The PLL can be much more stable now.

Koji

  2683   Thu Mar 18 19:00:04 2010 KojiSummaryElectronicsadvantege of our triple resonant EOM

Did you find what is the merit of their impedance matching technique?

Quote:

In this LVC meeting I discussed about triple resonant EOMs with Volker who was a main person of development of triple resonant EOMs at University of Florida.

Actually his EOM had been already installed at the sites. But the technique to make a triple resonance is different from ours.

They applied three electrodes onto a crystal instead of one as our EOM, and put three different frequencies on each electrode.

For our EOM, we put three frequencies on one electrode. You can see the difference in the attached figure. The left figure represents our EOM and the right is Volker's.

Then the question is; which can achieve better modulation efficiency ?

Volker and I talked about it and maybe found an answer,

 We believe our EOM can be potentially better because we use full length of the EO crystal.

This is based on the fact that the modulation depth is proportional to the length where a voltage is applied onto.

The people in University of Florida just used one of three separated parts of the crystal for each frequency.

 

  2682   Thu Mar 18 15:33:17 2010 kiwamuSummaryElectronicsadvantege of our triple resonant EOM

In this LVC meeting I discussed about triple resonant EOMs with Volker who was a main person of development of triple resonant EOMs at University of Florida.

Actually his EOM had been already installed at the sites. But the technique to make a triple resonance is different from ours.

They applied three electrodes onto a crystal instead of one as our EOM, and put three different frequencies on each electrode.

For our EOM, we put three frequencies on one electrode. You can see the difference in the attached figure. The left figure represents our EOM and the right is Volker's.

Then the question is; which can achieve better modulation efficiency ?

Volker and I talked about it and maybe found an answer,

 We believe our EOM can be potentially better because we use full length of the EO crystal.

This is based on the fact that the modulation depth is proportional to the length where a voltage is applied onto.

The people in University of Florida just used one of three separated parts of the crystal for each frequency.

Attachment 1: electrode.png
electrode.png
  2681   Thu Mar 18 13:40:35 2010 KojiUpdateABSLPLL reconstructed

We use the current PLL just now, but the renewal of the components are not immediate as it will take some time. Even so we need steady steps towards the better PLL. I appreciate your taking care of it.

Quote:

Quote:

Last night (Mar 17) I checked the PLL setup as Mott had some difficulty to get a clean lock of the PLL setting.

  • I firstly found that the NPRO beam is not going through the Faraday isolator well. This was fixed by aligning the steering mirrors before the Faraday.
     
  • The signal from the RF PD was send to the RF spectrum analyzer through a power splitter. This is a waist of the signal. It was replaced to a directional coupler.
  • Tee-ing the PZT feedback to the oscilloscope was producing the noise in the laser frequency. I put the oscilloscope to the 600Ohm output of the SR560, while connectiong the PZT output to the 50Ohm output.
  • In addition, 6dB+6dB attenuators have been added to the PZT feedback signal.

Now the beating signal is much cleaner and behave straight forward. I will add some numbers such as the PD DC output, RF levels, SR560 settings...

Now I am feeling that we definitely need the development of really clean PLL system as we use PLL everywhere! (i.e. wideband PD, nice electronics, summing amplifiers, stop poking SR560, customize/specialize PDH box, ...etc)

I also had noticed the progressive change of the aux NPRO alignment to the Farady.

I strongly agree about the need of a good and robust PLL.

By modifying the old PDH box (version 2008) eventually I was able to get a PLL robust enough for my purposes. At some point that wasn't good enough for me either.

I then decided to redisign it from scratch. I'm going to work on it. Also because of my other commitments, I'd need a few days/1 week for that. But I'd still like to take care of it. Is it more urgent than that?

 

  2680   Thu Mar 18 12:27:56 2010 AlbertoUpdateABSLPLL reconstructed

Quote:

Last night (Mar 17) I checked the PLL setup as Mott had some difficulty to get a clean lock of the PLL setting.

  • I firstly found that the NPRO beam is not going through the Faraday isolator well. This was fixed by aligning the steering mirrors before the Faraday.
     
  • The signal from the RF PD was send to the RF spectrum analyzer through a power splitter. This is a waist of the signal. It was replaced to a directional coupler.
  • Tee-ing the PZT feedback to the oscilloscope was producing the noise in the laser frequency. I put the oscilloscope to the 600Ohm output of the SR560, while connectiong the PZT output to the 50Ohm output.
  • In addition, 6dB+6dB attenuators have been added to the PZT feedback signal.

Now the beating signal is much cleaner and behave straight forward. I will add some numbers such as the PD DC output, RF levels, SR560 settings...

Now I am feeling that we definitely need the development of really clean PLL system as we use PLL everywhere! (i.e. wideband PD, nice electronics, summing amplifiers, stop poking SR560, customize/specialize PDH box, ...etc)

I also had noticed the progressive change of the aux NPRO alignment to the Farady.

I strongly agree about the need of a good and robust PLL.

By modifying the old PDH box (version 2008) eventually I was able to get a PLL robust enough for my purposes. At some point that wasn't good enough for me either.

I then decided to redisign it from scratch. I'm going to work on it. Also because of my other commitments, I'd need a few days/1 week for that. But I'd still like to take care of it. Is it more urgent than that?

  2679   Thu Mar 18 10:46:51 2010 KojiUpdateABSLPLL reconstructed

Last night (Mar 17) I checked the PLL setup as Mott have had some difficulty to get a clean lock of the PLL setting.

  • I firstly found that the NPRO beam is not going through the Faraday isolator well. This was fixed by aligning the steering mirrors before the Faraday.
     
  • The signal from the RF PD was send to the RF spectrum analyzer through a power splitter. This is a waist of the signal. It was replaced to a directional coupler.
     
  • Tee-ing the PZT feedback to the oscilloscope was producing the noise in the laser frequency. I put the oscilloscope to the 600Ohm output of the SR560, while connectiong the PZT output to the 50Ohm output.
     
  • In addition, 6dB+6dB attenuators have been added to the PZT feedback signal.

Now the beating signal is much cleaner and behave straight forward. I will add some numbers such as the PD DC output, RF levels, SR560 settings...

Now I am feeling that we definitely need the development of really clean PLL system as we use PLL everywhere! (i.e. wideband PD, nice electronics, summing amplifiers, stop poking SR560, customize/specialize PDH box, ...etc)

  2678   Thu Mar 18 08:49:51 2010 steveUpdatePEMGuralp2 centered again ?

 

Someone adjusted the Guralp2 mass position last night??

NO

Attachment 1: grlp2xadj.jpg
grlp2xadj.jpg
  2677   Tue Mar 16 09:37:30 2010 steveUpdateSUSeq 4.4 seen by oplevs and osems

The oplev plots clearly show the alignment effect of this eq.

Attachment 1: opleveq4.4.jpg
opleveq4.4.jpg
Attachment 2: eq4.4.jpg
eq4.4.jpg
Attachment 3: opleveq4.4d3.jpg
opleveq4.4d3.jpg
  2676   Tue Mar 16 05:10:39 2010 kojiUpdateEnvironmentEarthquake, Mag 4.4

Some of the suspensions got watchdog tripped -> enabled -> damped.

The MC mirrors got slightly misaligned.

 

Attachment 1: 100316_MC_SUS.png
100316_MC_SUS.png
  2675   Tue Mar 16 04:22:02 2010 JenneUpdateEnvironmentEarthquake, Mag 4.4

Earthquake Details

Magnitude 4.4
Date-Time
  • Tuesday, March 16, 2010 at 11:04:00 UTC
  • Tuesday, March 16, 2010 at 04:04:00 AM at epicenter
Location 33.998°N, 118.072°W
Depth 18 km (11.2 miles)
Region GREATER LOS ANGELES AREA, CALIFORNIA
Distances
  • 2 km (1 miles) ENE (58°) from Pico Rivera, CA
  • 4 km (3 miles) ESE (115°) from Montebello, CA
  • 4 km (3 miles) SSW (207°) from Whittier Narrows Rec. Area, CA
  • 9 km (5 miles) NE (39°) from Downey, CA
  • 9 km (6 miles) SSW (206°) from El Monte, CA
  • 17 km (11 miles) ESE (112°) from Los Angeles Civic Center, CA
Location Uncertainty horizontal +/- 0.4 km (0.2 miles); depth +/- 0.9 km (0.6 miles)
Parameters Nph=122, Dmin=5 km, Rmss=0.37 sec, Gp= 25°,
M-type=local magnitude (ML), Version=2
Source
Event ID

ci14601172

-----------------------------------------

MC1 and MC3 seem to have kept themselves together, but all the other optics' watchdogs tripped.

Attachment 1: EQ_16Mar2010.png
EQ_16Mar2010.png
  2674   Mon Mar 15 16:39:36 2010 steveUpdatePEMGuralp2 centered

Quote:

Untitled.png

 

Guralp 2 centered.The mass position offsets are: E-W 0.05V, N-S 0V, Z 0.4V

Guralp 1: E-W -0.1V, N-S -0.25V, Z 0V measured, not adjusted

The GUR2_X channel has an offset. See plot below when seismometers are disconnected. This offset has to be removed.

NOTE: this huddle is on bad-soft ground-lenoleum tile from prehistoric Flintstone age

 

Attachment 1: grlp2ffst.jpg
grlp2ffst.jpg
  2673   Mon Mar 15 09:43:47 2010 steveUpdatePEMmore sandblasting today

Do not open IFO vacuum envelope today! They are sandblasting again at CES

  2672   Sat Mar 13 22:16:03 2010 ranaUpdateWienerFilteringGuralp Huddle Test software

I used some recent better data to try for better Z subtraction.

Dmass helped me understand that sqrt(1-Coherence) is a good estimate of the theoretical best noise subtraction residual. This should be added to DTT. For reference the Jan statistic is the inverse of this.

This should get better once Steve centers the Guralps. 

Attachment 1: guralpznoise.png
guralpznoise.png
  2671   Sat Mar 13 21:30:22 2010 ranaUpdatePEMStyrofoam Cooler on the Seismos

Untitled.png

 

  2670   Fri Mar 12 17:08:22 2010 steveConfigurationVACbg-RGAscan at d18

RGA scan of rga-region only at day 18   This is the back ground of the rga with some calibration gas.

Attachment 1: bg-d18_20100312scan.png
bg-d18_20100312scan.png
  2669   Fri Mar 12 13:52:18 2010 ZachUpdateelogelog restarted

 The elog was down and I ran the restart script.

  2668   Thu Mar 11 17:51:04 2010 KojiUpdateSUSRecent status of SOSs

Jenne, Koji

Recent status of SOSs:

We completed one of the suspension (ITMY).

ITMX: 6 Magnets, standoffs, and guide rod glued / balance to be confirmed / needs to be baked
ITMY: 6 Magnets, standoffs, and guide rod glued / balance confirmed / needs to be baked
SRM: 6 Magnets, one standoff, and guide rod glued,  / waiting for the release from the gluing fixture.
PRM: one standoff, and guide rod glued / waiting for the magnet gluing. 

We think we solved all the problems for hanging the suspensions.

--- Magnet gluing fixture ---

  • There is the two kinds of fixtures. Neither does work propery in the original form!
  • The height of the side magnets should be finely adjusted by changing the teflon sheets beneath the optics in the fixture.
  • Be aware of the polarity of the fixture in terms of the side magnets
  • Wrongly glued magnets (and others) can be removed by a razor blade with some amount of acetone.
  • The pickle picker frequently knocks the magnets down during the release. Don't s be down in the dumps too much.

--- Suspending the mirror ---

  • The wire winches must be carefully attached to the suspension tower such that the wires are not streached during fastening the clamps.
  • There are a couple variations of the drawings for SOS. The one we have has #4-40 for the earthquake stops at the bottom.
    Zach and Mott made the EQ stops with the right size.
  2667   Thu Mar 11 15:24:51 2010 steveMetaphysicsEnvironmentIFO was well protected

Quote:

There is a planned power outage tomorrow, Saturday from 7am till midnight.

I vented all annulies and switched to ALL OFF configuration. The small region of the RGA is still under vacuum.

The vac-rack: gauges, c1vac1 and UPS turned off.

 It turns out that we perfecly timed the big one

In the process of finding the signal of the big chilean earthquake I just realized that we were all off

  2666   Wed Mar 10 15:02:58 2010 ranaUpdatePEMStyrofoam Cooler on the Seismos

This is the spectra and coherence from a quiet time last night. I've lowered the Guralp cal by a factor of 2 to account for the fact that the gain in the breakout box is actually 20 and not 10 as I previously said.

The AD620 stage in the front part has a gain of 10 and then there's a single-to-differential stage in the output which gives us a gain of 2. The DTT cal in counts is now 3.8e-9 (m/s)/count.

The second plot shows the Guralp and Ranger signals at the ADC input (converted from counts to Volts for usefulness). The thick grey line is the expected noise of the Guralp breakout box

(mainly the AD620) propagated to the ADC (via multiplication by 2). It looks like the preamp board should not be a problem as long as we can reach the AD620 limit.

So the excess noise in the Guralp is not the fault of the preamp, but more likely the mounting and insulation of the seismometers.

Attachment 1: Seismic-Ref_100310.png
Seismic-Ref_100310.png
Attachment 2: Untitled.gif
Untitled.gif
  2665   Tue Mar 9 12:06:53 2010 rana, JenneUpdatePEMStyrofoam Cooler on the Seismos

Looks like the GUR2_X signal is bad. Jenne says that we need to center it mechanically before the signals will become useful again. Maybe Steve will do this - instructions are in the manuel ?

  2664   Tue Mar 9 09:32:31 2010 KojiSummaryGeneralWideband measurement of Fast PZT response

I have measured a wideband response of the fast PZT in the LWE NPRO 700mW in the Alberto's setup.
This is a basic measurement to determine how much phase modulation we can obtain by actuating the fast PZT,
primarily for the green locking experiment.

RESULT

  • Above 200kHz, there are many resonances that screws up the phase.
     
  • Modulation of 0.1rad can be easily obtained even at 10MHz if the modulation frequency is scanned.
     
  • Change of the laser frequency in DC was observed depending on the modulation frequency.
    i.e. At the resonance the laser frequency escaped from the RF spectrum analyzer.
    This may induced by the heat dissipation in the PZT causing the temperature change of the crystal.
     
  • Some concerns: Is there any undesired AM by the PZT modulation?

---

METHOD

1. Locked the PLL of for the PSL-NPRO beating at 20MHz.

2. Added the modulation signal to the NPRO PZT input.
I used the output of the network analyzer sweeping from 100kHz to 1MHz.

3. Measured the transfer function from the modulation input to the PLL error signal.
The PLL error is sensitive to the phase fluctuation of the laser. Found that the first resonance is at 200kHz.
The TF is not valid below 3kHz where the PLL suppresses the modulation.

4. Single frequency modulation: Disconnected the PLL setup.
Plug Marconi into the fast PZT input and modulate it at various frequencies.
Observing with the RF spectrum analyzer, I could see strong modulation below 1MHz.
It turned out later that the TF measurement missed the narrow peaks of the resonances due to the poor freq resolution.

Also the modulation depth varies frequency by frequency because of the resonances.
Scanned the frequency to have local maximum of the modulation depth. Adjusted the
modulation amplitude such that the carrier is suppressed
(J0(m)=0 i.e. m~2.4). As I could not obtain
the carrier suppression at above 1MHz, the height of the carrier and the sidebands were measured.

The modulation frequency was swept from 100kHz to 10MHz.

5. Calibration. The TF measured has been calibrated using the modulation depth obtained at 100Hz,
where the resonance does not affect the response yet.

The responce of the PZT was ~10MHz/V below 30kHz. Looks not so strange although this valure is
little bit high from the spec (2MHz/V), and still higher than my previous experience at TAMA (5MHz/V).
Note that this calibration does not effect to the modulation depth of the single freq measurement as they are independent.

Attachment 1: PZT_response.png
PZT_response.png
  2663   Tue Mar 9 09:04:20 2010 steveUpdatePEMkeep vacuum chamber closed

They are sandblasting at CES: our particle counts are very high. DO NOT OPEN CHAMBER!

Attachment 1: sandblasting.jpg
sandblasting.jpg
  2662   Mon Mar 8 18:15:27 2010 ranaUpdatePEMStyrofoam Cooler on the Seismos

I put Jenne's cooler over the seismometers. Kiwamu put the copper foil wrapped lead brick on top of the cooler to hold it down. I also put another (unwrapped) lead brick on top of the Guralp cables outside of the cooler. Frank gave me a knife with which I cut a little escape hole in the bottom of the cooler lip for the Guralp cables to sneak out of.

Attachment 1: SeismometersHuddled.jpg
SeismometersHuddled.jpg
  2661   Sun Mar 7 23:05:39 2010 ranaUpdatePEMSeismic witnesses near MC1 tank moved

Quote:

Since we're going to open the MC1 tank tomorrow, I've moved the MC1 accelerometers and the Guralp over to underneath MC2 for the vent.  I'll reconnect them later.

 I've put both Guralps next to the Ranger and connected them to the breakout box. The data is now good.

I found that the Ranger was not centered and so it was stuck (someone kicked it in the last 2 weeks apparently). I recentered the mass according to the procedure in the manual. Its now moving freely.

In order to do a better huddle test, I increased the gain of the Ranger's SR560 preamp to 100 from 10 and put it on the low noise setting. I also enabled a 2x lowpass at 3 kHz for no good reason.

I couldn't find what the actual value of the gain of the Guralp breakout box is, but I assume its 10. With this assumption the calibrations are this:

Guralp: 800 V/(m/s)  *  10  (V/V)   *  16384  cts/V   =>    7.63e-9  (m/s)/count           (0.03 - 40 Hz)

Ranger:  345 V/(m/s)  * 100 (V/V)  *  16384 cts/V   =>     1.77e-9  (m/s)/count          (above 1Hz)

To account for the fact that I am not damping the Ranger with an external damping resistor, I have changed the calibration poles and zeros: in DTT we now use 2 poles @ 0 Hz and a complex pair at 1 Hz:

G = 1.77e-9

Poles = 0, 0

Zeros = 0.15 0.9887

I think that the Guralp gain is too high by a factor of 2. To really do this right, we should attach a known voltage to the input pins of the Guralp breakout and then read off the amount of counts.

Attachment 1: seis.png
seis.png
  2660   Sun Mar 7 07:01:21 2010 ranaUpdateWienerFilteringGuralp Huddle Test software

We need to do a new huddle test of the Guralps for the Wiener filtering paper. The last test had miserable results.

I tried to use recent data to do this, but it looks like we forgot to turn the Guralp box back on after the power outage or that they're far off center.

So instead I got data from after the previous power outage recovery.

I tried to use our usual Wiener filter method to subtract Guralp1-Z from Guralp2-Z, but that didn't work so well. It was very sensitive to the pre-weighting.

Instead I used the new .m file that Dmass wrote for subtracting the phase noise from his doubling noise MZ. That worked very well. It does all of the subtraction in the frequency domain and so doesn't have to worry about making a stable or causal filter. As you can see, it beats our weighted Wiener filter at all frequencies.

huddlez.png

The attached plot shows the Guralp spectra (red & green), the residual using time-domain Wiener filtering (black) and the Dmass f-domain code (yellow).

As soon as Jenne brings in her beer cooler, we're ready to redo the Huddle Test.

 

  2659   Fri Mar 5 18:04:56 2010 ranaUpdateSUSused OSEMs are magnetic

The OSEM LEDs and PDs from Honeywell have always had some ferromagnetic material in them. These are the same OSEMs we had since 2000.

You must be thinking of the really old 20th century plastic OSEMs.

  2658   Fri Mar 5 11:21:18 2010 steveUpdateSUSused OSEMs are magnetic

Quote:

Jenne and Koji

We successfully hung ITMX on the SOS. Side magnet is ~2mm off from the center of the OSEM. ITMX aligned using the QPD. The OSEMs changes the alignment. It looks that something magnetic is inside the OSEM PD or LED.

Reguled ITMY side magnet.

Cleaned up the lab for the safety inspection.

 The brand new OSEM LED and PD can be picked up with a weak magnet. These ferrous metals of LEDs and PDs will be magnetized by sitting in the sus next to the

magnets for years. I hanged optics with new OSEMs and never saw this effect before.

 

 

We have to demagnetize them.

  2657   Thu Mar 4 22:07:21 2010 ranaUpdate40m Upgrading11MHz PD not yet designed

Use 10 Ohms for the resistance - I have never seen a diode with 25 Ohms.

p.s. PDFs can be joined together using the joinPDF command or a few command line options of 'gs'.

  2656   Thu Mar 4 19:53:56 2010 AlbertoUpdate40m Upgrading11MHz PD designed adjusted for diode's resistance; 55 MHz RFPD designed
After reading this study done at LIGO MIT in 1998 I understood why it is difficult to define an effective impedance for a photodiode.

I read a few datasheets of the C30642GH photodiode that we're going to use for the 11 and 55 MHz. Considering the  values listed for the resistance and the capacitance in what they define "typical conditions" (that is, specific values of bias voltage and DC photocurrent) I fixed Rd=25Ohms and Cd=175pF.

Then I picked the tunable components in the circuit so that we could adjust for the variability of those parameters.

Finally with LISO I simulated transfer functions and noise curves for both the 11 and the 55MHz photodiodes.

I'm attaching the results and the LISO source files.

 

Attachment 1: rfpd55_Noise.pdf
rfpd55_Noise.pdf
Attachment 2: rfpd55_TF.pdf
rfpd55_TF.pdf
Attachment 3: rfpd11_v2_TF.pdf
rfpd11_v2_TF.pdf
Attachment 4: rfpd11_v2_Noise.pdf
rfpd11_v2_Noise.pdf
Attachment 5: rfpd11_v2.fil
Attachment 6: rfpd55.fil
  2655   Thu Mar 4 08:43:35 2010 AlbertoUpdate40m Upgrading11 MHz RFPD elctronics

Quote:
** Please add LISO file w/ component values.

oops, forgotten the third attachment...

here it is

Attachment 1: rfpd11_v2.fil
# Resonant RF diode front end
#
#		  gnd
#		  |
#		  Cw2
#		  |
#		  n23
#		  |
#		  Lw2
#		  |
... 60 more lines ...
  2654   Thu Mar 4 02:25:14 2010 JenneUpdateCOCFurther details on the magnet story, and SRM guiderod glued

[Koji, Jenne]

First, the easy story:  SRM got it's guiderod & standoff glued on this evening.  It will be ready for magnets (assuming everything is sorted out....see below) as early as tomorrow.  We can also begin to glue PRM guiderods as early as tomorrow.

The magnet story is not as short.....

Problem: ITMX and ITMY's side magnets are not glued in the correct places along the z-axis of the optic (z-axis as in beam propagation direction). 

ITMX (as reported the other day) has the side magnet placement off by ~2mm.  ITMX side was glued using the magnet fixture from MIT and the teflon pads that Kiwamu and I improvised.

It was determined that the improvised teflon pads were too thin (maybe about 1m thick), so I took those out, and replaced them with the teflon pads stolen from the 40m's magnet gluing fixture.   (The teflon pad from the MIT fixture and the ones from the MIT fixture are the same within my measuring ability using a flat surface and feeling for a step between them.  I haven't yet measured with calipers the MIT pad thickness).  The pads from the 40m fixture, which were used in the MIT fixture to glue ITMY side last night were measured to be ~1.7mm thick.

Today when Koji hung ITMY, he discovered that the side magnet is off by ~1mm.  This improvement is consistent with the switching of the teflon pads to the ones from the 40m fixture.

We compared the 40m fixture with the one from MIT, and it looks like the distance from the edge of where the optic should sit to the center of the hole for the side magnet is different by ~1.1mm.  This explains the remaining ~1mm that ITMY is off by. 

We should put the teflon pads back into the 40m fixture, and only use that one from now on, unless we find an easy way to make thicker teflon pads for the fixture we received from MIT.  (The pads that are in there are about the maximum thickness that will fit).  I'm going to use my thickness measurements of SRM (taken in the process of gluing the guiderods) to see what thickness of pads / what fixture we want to actually use, but I'm sure that the fixture we found in the 40m is correct.  We can't use this fixture however, until we get some clean 1/4-28 screws.  I've emailed Steve and Bob, so hopefully they'll have something for us by ~lunchtime tomorrow. 

The ITMX side magnet is so far off in the Z-direction that we'll have to remove it and reglue it in the correct position in order for the shadow sensor to do anything.  For ITMY, we'll check it out tomorrow, whether the magnet is in the LED beam at all or not.  If it's not blocking the LED beam enough, we'll have to remove and reglue it too. 

Why someone made 2 almost identical fixtures, with a 1mm height difference and different threads for the set screws, I don't know.  But I don't think whoever that person was can be my friend this week. 

  2653   Wed Mar 3 18:32:25 2010 AlbertoUpdate40m Upgrading11 MHz RFPD elctronics
** Please add LISO file w/ component values.
 
I designed the circuit for one of the 11 MHz photodiodes that we're going to install in the 40m Upgrade.

This is a simple representation of the schematic:

          gnd
#          |
#          Cw2
#          |
#          n23
#          |
#          Lw2
#          |
#           n22
#          |
#          Rw2                
#                 |                   |\            
#           n2- - - C2 - n3 -  - -  - |  \          
#            |    |      |   |        |4106>-- n5 - Rs -- no
# iinput    Rd   L1     L2 R24    n6- |  /     |           |
#      nin - |    |      |   |    |   |/       |         Rload    
#           Cd   n7     R22 gnd   |            |           |          
#            |    |      |        | - - - R8 - -          gnd              
#           gnd  R1     gnd      R7 
#                 |               |
#         gnd               gnd
#                 
#
#

I chose the values of the components in a realistic way, that is using part available from Coilcraft or Digikey.

Using LISO I simulated the Tranfer Function and the noise of the circuit.

I'm attaching the results.

I'll post the 55MHz rfpd later.

Attachment 1: rfpd11_v2_TF.pdf
rfpd11_v2_TF.pdf
Attachment 2: rfpd11_v2_Noise.pdf
rfpd11_v2_Noise.pdf
  2652   Wed Mar 3 02:57:29 2010 JenneUpdateSUSITMX hung

Quote:

Jenne and Koji

We successfully hung ITMX on the SOS. Side magnet is ~2mm off from the center of the OSEM.

Reguled ITMY side magnet.

Cleaned up the lab for the safety inspection.

 Some details on the side magnet situation from today: 

To glue the magnets+dumbbells to the optics, we use the magnet-dumbbell gluing fixture.  This fixture is supposed to have teflon 'pads' for the optic to sit on while you align it in the fixture, however the fixture which we received from MIT (it's Betsy's....but it came via MIT) only had one of the 4 teflon pads. 

Kiwamu and I decided (last week, when we first glued ITMX's magnets) that it would be bad news to let the AR face of the optic sit on bare metal, so we fashioned up some teflon pads using stock in a cabinet down the Yarm.  We were focused on thinking about the face magnets, and didn't think about how the thickness of the teflon affected the placement of the side magnet.  We chose some teflon that was too thin by ~1mm, so the optic sat too low in the fixture, resulting in the side magnet being glued too close to the HR side of the optic (this is all along the Z - axis, where Z is the direction of beam propagation). 

Why it ended up being 2mm off instead of only 1mm I don't really have an explanation for, other than perhaps tightening the set screws to hold the optic (by the barrel) in the fixture pushes the optic up.  I observed this happening when I didn't put any effort into keeping the optic flat on the teflon pads, but I thought that I made sure the optic was seated nicely in the fixture before starting to glue.  When I glued the new ITMY side magnet tonight I tried to make sure that the optic was seated nicely in the fixture.  We'll see what happens.

Before gluing the new ITMY side magnet (and now it's set for all future magnet gluings....), I found 4 teflon pads of all the correct thickness.  It turns out that we have a magnet gluing fixture of our own, which I found in the cabinets in the clean room.  This fixture had all 4 teflon pads, so I stole them and put them into the one that we're using for this round of upgrade / suspension hangings.  The height of all future side magnets should be correct. The thickness of the pads in the 'spare' fixture matched the one which came with the fixture from MIT as closely as I could feel by putting them on the same flat surface next to each other and feeling if there was a step. 

A side note about this magnet gluing fixture that I found:  It has the word "TOP" etched into it, to prevent exactly my problem with the ITMY side magnets in the first place.  Unfortunately the threads for the set screws which hold the optic are shot (or something is funny with them), so we can't just use this fixture. 

Gluing notes regarding the standoffs and guiderods:

There's more glue than I'd like on the guiderods / standoff for ITMX.  The glue was starting to get a little tacky when I glued the standoff in place after we balanced the optic, so it was hard to get it in the right place.  I'm confident we have a good epoxy contact, and we don't have much glue that I think it'll be a big problem.  Certainly I'll be a lot better at manuvering my glue-stick a.k.a skinny piece of wire around the suspension tower to get to the standoff for the rest of the optics that we're hanging, and I won't have glued something like ITMY side magnet immediately beforehand, which took enough time that the glue started to get tacky (not very tacky, just barely noticeably tacky).

I'd say that most gluing activities should be completed within ~10-15min of mixing the glue, after spending ~2min stirring to make sure it's nice and uniform.  It doesn't dry fast enough to be a huge rush, but you should get right on the gluing once the epoxy has been mixed. 

  2651   Tue Mar 2 23:11:43 2010 KojiUpdateSUSITMX hung

Jenne and Koji

We successfully hung ITMX on the SOS. Side magnet is ~2mm off from the center of the OSEM. ITMX aligned using the QPD. The OSEMs changes the alignment. It looks that something magnetic is inside the OSEM PD or LED.

Reguled ITMY side magnet.

Cleaned up the lab for the safety inspection.

  2650   Tue Mar 2 12:20:54 2010 kiwamuUpdatePSLstray beam

In order to block stray beams, I have put some beam dumps and razor blades on the PSL  table.

There were three undesired spots in total. I found two spots on the south side door of the PSL room, close to Mach-Zehnder.

Another spots was on the middle of the north door. Now they all are blocked successfully.

  2649   Mon Mar 1 22:38:12 2010 ranaUpdateComputersRC sensitivity to RIN

The overnight triangle wave I ran on the AOM drive turns out to have produced no signal in the FAST feedback to the PZT.

The input power to the cavity was ~10 mW (I'm totally guessing). The peak-peak amplitude of the triangle wave was 50% of the total power.

The spectral density of the fast signal at the fundamental frequency (~7.9 mHz) is ~0.08 V/rHz. The FAST calibration is ~5 MHz/V. So, since we

see no signal, we can place an upper limit on the amount of frequency shift = (5 MHz/V) * (0.08 V/rHz) * sqrt(0.0001 Hz) = 4 kHz.

Roughly this means that the RIN -> Hz coefficient must be less than 4 kHz / 5 mW or ~ 1 Hz/uW.

For comparison, the paper on reference cavities by the Hansch group lists a coefficient of ~50 Hz/uW. However, they have a finesse of 400000

while we only have a finesse of 8000-10000. So our null result means that our RC mirrors' absorption is perhaps less than theirs. Another possibility

is that their coating design has a higher thermo-optic coefficient. This is possible, since they probably have much lower transmission mirrors. It would be

interesting to know how the DC thermo-optic coefficient scales with transmission for the standard HR coating designs.

Attachment 1: Untitled.png
Untitled.png
  2648   Mon Mar 1 11:21:50 2010 steveConfigurationVACvacuum contoller is back to normal

Quote:

There is a planned power outage tomorrow, Saturday from 7am till midnight.

I vented all annulies and switched to ALL OFF configuration. The small region of the RGA is still under vacuum.

The vac-rack: gauges, c1vac1 and UPS turned off.

 Vac- rack is powered back up.  UPS first, than all other power switches from top to bottom of the rack, except Maglev

Manually started one by one TP2 and TP3  to accelerate to 50 KRPM

Brought up vac.control screen on lap-top at /cvs/cds/caltech/medm/c0/ve/VacControl_BAK.adl

V5 and VM3 were opened so TP3 can pump on the RGA

V4 was opened so TP2 can pump on the Maglev-TP1. The Maglev power was  turned on and started acceleration.

The vac control screen positions indicators were checked for true position and annulies vent valves were opened.

RGA manual on/off switch was turned at the top of the RGA-head. Ubuntu copmuter was started at cc4 1.1e-6 Torr

The RGA communication was started with:  ssh c0rga from control room

The rga-script was started ./RGAset.py     This script turns on the filament, rs-232 and scan parameters etc

 

Vac -configuration: IFO-P1 at atm, RGA is pumped and running in background mode, all annulos at atm

 

  2647   Mon Mar 1 08:49:37 2010 steveBureaucracySAFETYsafety audit

The 40m safety audit will be at Wednesday afternoon, March 3

Please participate getting our lab to inspection grade safety level.

  2646   Sun Feb 28 23:47:52 2010 ranaUpdateComputersrosalba

Since Rosalba wanted to update ~500 packages, I let it do it. This, of course, stopped the X server from running. I downloaded and installed the newest Nvidia driver and its mostly OK.

The main problem with the auto-update on our workstations is that we've updated some packages by hand; i.e. not using the standard CentOS yum. So that means that the auto-update doesn't work right. From now on, if you want to install a fancier package than what CentOS distributes, you should commit to handle the system maintenance for these workstations for the future. Its not that we can't have new programs, we just have to pay the price.

 

At 23:45 PST, I also started a slow triangle wave on the AOM drive amplitude. This is to see if there's a response in the FSS-FAST which might imply a coupling from intensity noise to frequency noise via absorbed power and the dn/dT effect in the coatings.

Its a 93 second period triangle modulating the RC power from 100% down to 50%.

  2645   Sun Feb 28 16:45:05 2010 ranaSummaryGeneralPower ON Recovery
  1. Turned ON the RAID above linux1.
  2. Hooked up a monitor and keyboard and then turned ON linux1.
  3. After linux1 booted, turned ON nodus - then restarted apache and elog on it using the wiki instructions.
  4. Turned on all of the control room workstatiions, tuned Pandora to Johnny Cash, started the auto package updater on Rosalba (517 packages).
  5. Started the startStrip script on op540m.
  6. turned on RAID for frames - wait for it to say 'SATA', then turn on daqctrl and then fb40m and then daqawg and then dcuepics
  7. turned on all the crates for FEs, Sorensens, Kepcos for LSC, op340m, mafalda was already on
  8. fb40m again doesn't mount the RAID again!
  9. I turned on fb40m2 and that fixes the problem. The fb40m /etc/vfstab points to 198.168.1.2, not the JetStor IP address.
  10. I plugged in the Video Switch - its power cord was disconnected.
  11. FEs still timing out saying 'no response from EPICS', but Alberto is now here.

Sun Feb 28 18:23:09 2010

Hi. This is Alberto. Its Sun Feb 28 19:23:09 2010

  1. Turned on c1dcuepics, c0daqctrl and c0daqawg. c0daqawg had a "bad"status on the daqdetail medm screen. The FEs still don't come up.
  2. Rebooted c1dcuepics and power cycled c0daqctrl and c0daqawg. The problem is still there.
  3. Turned on c1omc. Problem solved.
  4. Rebooted c1dcuepics and power cycled c0daqctrl and c0daqawg. c0daqawg now good. The FE are coming up.
  5. Plugged in the laser for ETMY's oplev
  6. Turned on the laser of ETMX's oplev from its key.

 Monday, March 1, 9:00 2010 Steve turns on PSL-REF cavity ion pump HV at 1Y1

  2644   Fri Feb 26 15:32:13 2010 steveConfigurationVACpreparation for power outage: vacuum all off

There is a planned power outage tomorrow, Saturday from 7am till midnight.

I vented all annulies and switched to ALL OFF configuration. The small region of the RGA is still under vacuum.

The vac-rack: gauges, c1vac1 and UPS turned off.

Attachment 1: ventd3.jpg
ventd3.jpg
ELOG V3.1.3-